OpenStudy (anonymous):

Write the equation of the line below in slope intercept form. CAN SOMEBODY PLEASE EXPLAIN TO ME HOW TO DO THIS I AM SO CONFUSED WILL GIVE MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

i can help, what equation?

OpenStudy (anonymous):

^^^^^^^^^^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

I do not want the answer i just want to know how to write the equation for a line in slope intercept form. @DuckDynastyfan923 and @iGreen.

OpenStudy (anonymous):

oh, ok. slope intercept form is y=mx+b.

OpenStudy (anonymous):

b is the y intercept, and m is slope.

OpenStudy (anonymous):

You have to rearrange it to slope-intercept form.. y = mx + b Let's say we have: 2x + 5y = 9 To re-arrange, we want 'y' by itself on one side. So first we subtract 2x to both sides: 5y = -2x + 9 Now we divide 5 to both sides: y = -2/5x + 9/5 So this is 2x + 5y = 9 in slope intercept form.

OpenStudy (anonymous):

Do you understand? The whole point is to get 'y' by itself.

OpenStudy (anonymous):

but I have to write an equation for the line and I do not have any numbers I just have a line on a graph. so how would I do that? @iGreen.

OpenStudy (anonymous):

find two points on the line, and find slope using this equation- say your two points were (3,4) and (4,3). put them into the equation, and you have  then just simplify.